6. Two impedances = ( 4 + j7 )Ω and z2 = ( 4 - 3j )Ω are connected in parallel determine:


a. the current flowing

b. its phase relative to the 120V supply voltage.


Total impedance for a parallel circuit is given by 1/Z = 1/Z + 1/Z2 and the current is given by I = V/Z





1
Expert's answer
2021-03-26T15:07:31-0400

"\frac{1}{Z} = \frac{1}{Z_1} + \frac{1}{Z_2}"


"\frac{1}{Z} = \frac{1}{4+j7} + \frac{1}{4-3j}"


"\frac{ 1}{Z} = \frac{4-3j + 4+7j}{(4+7j)(4-3j)}"


"\frac{1}{Z} = 0.22+0.012j"



V = 120 + 0j

I = "\frac{V}{Z}" = V × "\frac{1}{Z}"


I = (120+0j) × (0.22 + 0.012j)


I = 26.4 + 1.44j